OpenStudy (anonymous):

Part A: The area of a square is (16x2 − 8x + 1) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. Part B: The area of a rectangle is (81x2 − 4y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work.

OpenStudy (anonymous):

Part A By the identity \[\Large (a ^{2} \pm 2ab+b ^{2})=(a \pm b)^{2}\] \[\Large(16x^2 − 8x + 1)=(4x-1)^{2}\] so the side of the square is 4x-1 Part B By the identity \[\Large(a ^{2} - b ^{2})=(a-b)(a+b)\] \[\Large(81x^2 − 4y^2)=(9x-2y)(9x+2y)\] then the sides of the rectangle are 9x-2y and 9x+2y

OpenStudy (jtvatsim):

Does that help @SOAD_Fan ? :)

OpenStudy (anonymous):

Ja :)

OpenStudy (jtvatsim):

Awesome! Just wanted to make sure. @Natriumhydrid did a good job. :)

OpenStudy (jadedinsomniac):

SOAD fan huh what your fav album